### v3

@@ Appendix F General Convex Functions under Heavy-Tailed Noise
In this section, we provide the full version of Theorem 5.2 by using Lemma E.2,
Theorem F.1.
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\log T+\frac{\eta^{p}(M^{p}+\sigma^{p})}{\eta_{*}}\log T\right)\right$.
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}\log T}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}\log T}{T^{1-\frac{1}{p}}}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\log T\right)\right$.
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}(\log T)^{\frac{2}{p}-1}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}(\log T)^{\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{T^{2}}\left(\sum_{t=1}^{T}\frac{1}{\eta_{t}}\right)+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{(T-t+1)^{2}}\left(\sum_{s=t}^{T}\frac{1}{\eta_{s}}\right)$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{\eta_{*}T^{2-\frac{2}{p}}}+\frac{5D_{\psi}(x,x^{1})}{3\eta T^{1-\frac{1}{p}}}$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{T}\left(\frac{T^{\frac{2-p}{p}}}{\eta_{*}}\lor\frac{T^{\frac{1}{p}}}{\eta}\right)+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\left(\frac{\eta_{*}}{T^{\frac{2-p}{p}}}\land\frac{\eta}{T^{\frac{1}{p}}}\right)^{\frac{2p-2}{2-p}}\sum_{t=1}^{T}\frac{1}{T-t+1}$
@@ F.1 Optimal Rate under Heavy-Tailed Noise
The full version of Theorem 5.3 is stated as follows.
Theorem F.2.
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\right)+\frac{1}{T^{1-\frac{1}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta}+\eta^{p-1}(M^{p}+\sigma^{p})\right)\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{2D_{\psi}(x,x^{1})}{\eta_{*}T^{2-\frac{2}{p}}}+\frac{2D_{\psi}(x,x^{1})}{\eta T^{1-\frac{1}{p}}}+\frac{(2-p)^{2}2^{\frac{5p-6}{2-p}}L^{\frac{p}{2-p}}\eta_{*}^{\frac{2p-2}{2-p}}}{p(p-1)T^{2-\frac{2}{p}}}+\frac{4^{p-\frac{1}{2}}(M^{p}+\sigma^{p})\eta^{p-1}}{p(p-1)T^{1-\frac{1}{p}}}$
@@ Appendix G Unified Theoretical Analysis under Sub-Weibull Noise
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x^{*}+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
@@
$\mathbb{\mathbb{E}}\left[\left|\Xi_{t}\right|^{k}\mid\mathcal{F}^{t}\right]\leq e\Lambda_{t}^{k}\left(\frac{2k}{p}\right)^{\frac{2k}{p}}\Rightarrow\left\|\Xi_{t}\right\|_{k}\leq\Lambda_{t}e^{\frac{1}{k}}\left(\frac{2k}{p}\right)^{\frac{2}{p}},$
where $\left\|\cdot\right\|_{k}\coloneqq\left(\mathbb{\mathbb{E}}\left[\left|\cdot\right|^{k}\right]\right)^{\frac{1}{k}}$. Hence, for any $k\geq 1$,
@@
$\mathbb{\mathbb{E}}\left[R_{t}\mid\mathcal{F}^{t-1}\right]=R_{t-1}\mathbb{\mathbb{E}}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle-6w^{2}\sigma^{2}\left\|Z^{t}\right\|^{2}-2^{q-1}w^{q}\sigma^{q}\left\|Z^{t}\right\|^{q}\right)\mid\mathcal{F}^{t-1}\right]\overset{(a)}{\leq}R_{t-1},$
where $(a)$ is by applying Lemma 6.1 to $\mathbb{\mathbb{E}}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle\right)\mid\mathcal{F}^{t-1}\right]$. Then, we consider a stopping time
@@
$\Pr\left[\exists s\in\left[T\right],R_{s}>\frac{4}{\delta}\right]$ $=\Pr\left[\tau\leq T\right]=\mathbb{\mathbb{E}}\left[\mathds{1}\left[\tau\leq T\right]\right]<\frac{\delta}{4}\mathbb{\mathbb{E}}\left[R_{\tau}\mathds{1}\left[\tau\leq T\right]\right]$
$=\frac{\delta}{4}\mathbb{\mathbb{E}}\left[R_{\tau\land T}\mathds{1}\left[\tau\leq T\right]\right]\leq\frac{\delta}{4}\mathbb{\mathbb{E}}\left[R_{\tau\land T}\right]\overset{(b)}{\leq}\frac{\delta R_{0}}{4}=\frac{\delta}{4},$ (75)
@@
$\mathbb{\mathbb{E}}\left[R_{t}\mid\mathcal{F}^{t-1}\right]=R_{t-1}\mathbb{\mathbb{E}}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle-6w^{2}\sigma^{2}\left\|Z^{t}\right\|^{2}\right)\mid\mathcal{F}^{t-1}\right]\overset{(c)}{\leq}R_{t-1},$
@@ Appendix H General Convex Functions under Sub-Weibull Noise
$F(x^{T+1})-F(x^{*})\leq\frac{2D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (80)
@@ H.1 Optimal Rate under Sub-Weibull Noise
$F(x^{T+1})-F(x^{*})\leq\frac{2D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (82)

### v4

@@ Appendix F General Convex Functions under Heavy-Tailed Noise
In this section, we prove the full version of Theorem 5.2 by using Lemma E.2,
Theorem F.1 (Full version of Theorem 5.2).
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq$ $O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\log T+\frac{\eta^{p}(M^{p}+\sigma^{p})}{\eta_{*}}\log T\right)\right$.
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}\log T}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}\log T}{T^{1-\frac{1}{p}}}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq$ $O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\log T\right)\right$.
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}(\log T)^{\frac{2}{p}-1}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}(\log T)^{\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{T^{2}}\left(\sum_{t=1}^{T}\frac{1}{\eta_{t}}\right)+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{(T-t+1)^{2}}\left(\sum_{s=t}^{T}\frac{1}{\eta_{s}}\right)$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{\eta_{*}T^{2-\frac{2}{p}}}+\frac{5D_{\psi}(x,x^{1})}{3\eta T^{1-\frac{1}{p}}}$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{T}\left(\frac{T^{\frac{2-p}{p}}}{\eta_{*}}\lor\frac{T^{\frac{1}{p}}}{\eta}\right)+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\left(\frac{\eta_{*}}{T^{\frac{2-p}{p}}}\land\frac{\eta}{T^{\frac{1}{p}}}\right)^{\frac{2p-2}{2-p}}\sum_{t=1}^{T}\frac{1}{T-t+1}$
@@ F.1 Optimal Rate under Heavy-Tailed Noise
The full version of Theorem 5.3 is stated below.
Theorem F.2 (Full version of Theorem 5.3).
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\right)+\frac{1}{T^{1-\frac{1}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta}+\eta^{p-1}(M^{p}+\sigma^{p})\right)\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{2D_{\psi}(x,x^{1})}{\eta_{*}T^{2-\frac{2}{p}}}+\frac{2D_{\psi}(x,x^{1})}{\eta T^{1-\frac{1}{p}}}+\frac{(2-p)^{2}2^{\frac{5p-6}{2-p}}L^{\frac{p}{2-p}}\eta_{*}^{\frac{2p-2}{2-p}}}{p(p-1)T^{2-\frac{2}{p}}}+\frac{4^{p-\frac{1}{2}}(M^{p}+\sigma^{p})\eta^{p-1}}{p(p-1)T^{1-\frac{1}{p}}}$
@@ Appendix G Unified Theoretical Analysis under Sub-Weibull Noise
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{E}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x^{*}+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
@@
$\mathbb{E}\left[\left|\Xi_{t}\right|^{k}\mid\mathcal{F}^{t}\right]\leq e\Lambda_{t}^{k}\left(\frac{2k}{p}\right)^{\frac{2k}{p}}\Rightarrow\left\|\Xi_{t}\right\|_{k}\leq\Lambda_{t}e^{\frac{1}{k}}\left(\frac{2k}{p}\right)^{\frac{2}{p}},$
where $\left\|\cdot\right\|_{k}\coloneqq\left(\mathbb{E}\left[\left|\cdot\right|^{k}\right]\right)^{\frac{1}{k}}$. Hence, for any $k\geq 1$,
@@
$\mathbb{E}\left[R_{t}\mid\mathcal{F}^{t-1}\right]=R_{t-1}\mathbb{E}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle-6w^{2}\sigma^{2}\left\|Z^{t}\right\|^{2}-2^{q-1}w^{q}\sigma^{q}\left\|Z^{t}\right\|^{q}\right)\mid\mathcal{F}^{t-1}\right]\overset{(a)}{\leq}R_{t-1},$
where $(a)$ is by applying Lemma 6.1 to $\mathbb{E}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle\right)\mid\mathcal{F}^{t-1}\right]$. Then, we consider a stopping time
@@
$\Pr\left[\exists s\in\left[T\right],R_{s}>\frac{4}{\delta}\right]$ $=\Pr\left[\tau\leq T\right]=\mathbb{E}\left[\mathds{1}\left[\tau\leq T\right]\right]<\frac{\delta}{4}\mathbb{E}\left[R_{\tau}\mathds{1}\left[\tau\leq T\right]\right]$
$=\frac{\delta}{4}\mathbb{E}\left[R_{\tau\land T}\mathds{1}\left[\tau\leq T\right]\right]\leq\frac{\delta}{4}\mathbb{E}\left[R_{\tau\land T}\right]\overset{(b)}{\leq}\frac{\delta R_{0}}{4}=\frac{\delta}{4},$ (78)
@@
$\mathbb{E}\left[R_{t}\mid\mathcal{F}^{t-1}\right]=R_{t-1}\mathbb{E}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle-6w^{2}\sigma^{2}\left\|Z^{t}\right\|^{2}\right)\mid\mathcal{F}^{t-1}\right]\overset{(c)}{\leq}R_{t-1},$
@@ Appendix H General Convex Functions under Sub-Weibull Noise
$F(x^{T+1})-F(x^{*})\leq\frac{2D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$
@@ H.1 Optimal Rate under Sub-Weibull Noise
$F(x^{T+1})-F(x^{*})\leq\frac{2D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$